Evaluate the following Integrals

3)  ʃ-11 (x + 1/2) √(x2+x+1)dx

4)  ʃ (x+2) / √(x+3) dx



1
Expert's answer
2021-09-14T06:05:55-0400


3)


"\int_{-1}^1 (x +\frac{1}{2}) \sqrt{(x^2+x+1)}dx"


Solution

"\int (x +\frac{1}{2}) \sqrt{(x^2+x+1)}dx"


Substitute. "u=x^2+x+1"

"\frac{du}{dx}=2x+1 \to dx=\frac{1}{2x+1}du:"


"=\frac{1}{2} \int \sqrt{u}du"


Now solving:


"=\int \sqrt{u}du"


Apply power rule


"=\int u^ndu=\frac{u^{n+1}}{n+1}" with "n=\frac{1}{2}"


"=\frac{2u^{\frac{3}{2}}}{3}"


Plug in solved integrals


"=\frac{1}{2} \int \sqrt{u}du"


"=\frac{u^{\frac{3}{2}}}{3}"


Undo substitution "u=x^2+x+1:"


"=\frac{(x^2+x+1)^{\frac{3}{2}}}{3}+C"


Inserting the bounds

"=[\frac{(1^2+1+1)^{\frac{3}{2}}}{3}+C]
-[\frac{((-1)^2+(-1)+1)^{\frac{3}{2}}}{3}+C]"


"\sqrt{3}-\frac{1}{3}=1.398717474235544"






4)

"\smallint
\frac{(x+2) }{\sqrt{(x+3)}} dx"


Solution

"\smallint
\frac{x+2}{\sqrt{x+3}} dx"

Substitute "u=x+3"

"\frac{du}{dx}=1 \to dx=du"


"=\smallint\frac{u-1}{\sqrt{u}}du"


Expand


"=\smallint[{\sqrt{u}}-\frac{1}{\sqrt{u}}]du"


Apply linearity


"=\smallint{\sqrt{u}}du
-
\smallint
\frac{1}{\sqrt{u}}du"


Now solving


"=\smallint \sqrt{u} du"


Apply power rule:


"=\int u^ndu=\frac{u^{n+1}}{n+1}" with "n=\frac{1}{2}"


"=\frac{2u^{\frac{3}{2}}}{3}"


Now solving


"\int\frac{1}{\sqrt{u}}du"


Apply power rule with "n=-\frac{1}{2}:"


"=2\sqrt{u}"


Plug in solved integrals

"=\smallint{\sqrt{u}}du
-
\smallint
\frac{1}{\sqrt{u}}du"


"=\frac{2u^{\frac{3}{2}}}{3}-2\sqrt{u}"


Undo substitution "u=x+3:"


"=\frac{2(x+3)^{\frac{3}{2}}}{3}-2\sqrt{x+3}"


"=\frac{2(x+3)^{\frac{3}{2}}}{3}-2\sqrt{x+3} +C"



"=2[\frac{2(x+3)^{\frac{3}{2}}}{3}-\sqrt{x+3} ]+C"


"=\frac{2x(x+3)^{}}{3}+C"